OpenStudy (anonymous):

Which is the exact value of the expression the square root of 96. + the square root of 150. − the square root of 54.? A:6the square root of 54. B:54the square root of 6. C:6the square root of 6. D:6the square root of 32.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

\[\sqrt{96}=4\sqrt{6} ,\sqrt{150}=5\sqrt{6} , \sqrt{54}=3\sqrt{6} , 4+5-3=6\sqrt{6}\]

OpenStudy (anonymous):

they are the same roots of 6,so u just copy the square root of 6 and plus the 4 and 5 minus 3...do u get it??